There is a jar that is 74 cm tall and has a diameter of 29.6 cm, the jar stands vertically and is half loaded up with water (ρwater = 1 g/cm3). Discover the gauge pressure close to the base of the jar. What is the force on the jar's base because of the water?

P1 = Unit in kPa

F1 = Unit in lb

Presently, some oil with ρoil = 0.8 g/cm3 is poured on head of the water so the total level of two liquids together is h2 = 57.7 cm (the liquids don't blend and don't get compacted). What is the gauge pressure and the force on the base of the jar for this situation?

P2 = Unit kPa

F2 = Unit lb
Answer #1

1) So gauge pressure in the jar coloumn is

P =\rho_w gh

Here rho is water density = 1gm/cm^3 = 1000kg/m^3

g is gravitational acceleration = 9.8 m/s^2

h = 74cm/2 = 37 cm = 0.37 m

So we get,

P =3626 Pa = 3.63 kPa

So Force on the base of jar

F= P.A= P.(\pi r^2)= 3.63 kPa(\pi (\frac{0.296}{2})^2)= 249.7 Nin lb unit

F= \frac{ 249.7 N}{4.45} = 56.11 lb

2) So oil density is lower than water that's why it floats on water and oil height is then

h2 = 57.7 - 37 = 20.7 cm= 0.207 m

So, gauge pressure is

P_{oil}= \rho_{oil} gh_2 = 800kg/m^3 (9.8)(0.207)= 1622 Pa = 1.62 kPaSo total gauge pressure.

P_2= 3.63+1.62 kPa=5.25 kPa

So net force on Bottom is

F_2= P_2 A= 81.15 lb